H2O + V VO^3+ + 2H+ + 5e-
2e- + 2H+ + NO3- NO2 - + H2O
1
Expert's answer
2017-05-09T02:16:09-0400
H2O + V = VO3+ + 2H+ + 5ē; [x2] oxidation; V – reducing agent
2ē + 2H+ + NO3- = NO2- + H2O; [x5] reducing; NO3-; NO3- - oxidizing agent
2V + 6H+ + 5NO3- = 3H2O + 2VO3+ + 5NO2- .
Answer:
2V + 6H+ + 5NO3- = 3H2O + 2VO3+ + 5NO2.